# Exploring the mysterious high-ionization source powering [NeĀ V] in high- analog SBS0335-052Ā E with JWST/MIRI

### v1

@@ 3.4 Photoionization and shock models
SSP models with and without the ionizing output from X-ray binaries (XRBs) from Garofali etĀ al. (2024; G24 hereafter);
@@
Intermediate-massive black hole (IMBH) models from Richardson etĀ al. (2022; R22 hereafter);
@@
Shock models from Flury etĀ al. (2024; F24 hereafter) and the Mexican Million Models Shock Database222http://3mdb.astro.unam.mx:3686/, described in Alarie & Morisset (2019; AM19 hereafter).
For all these models, we considered only the two lowest metallicity grids available (i.e., 5% and 10% Zā models for G24, R22, F24, and 2% and 10% Zā for MP23), in accordance with SBSā0335-052Ā Eās metallicity range (Papaderos etĀ al. 2006; Nakajima etĀ al. 2024). Also, we took into account only bursts of star-formation with ages below 100Ā Myr (upper limit on age of SBS 0335-052Ā E) to match SBS 0335-052Ā E properties (Izotov etĀ al. 1997; Wofford etĀ al. 2021; Berg etĀ al. 2022; Mingozzi etĀ al. 2024, Martinez et al. in prep.). For all the grids, we consider the ionization parameter log() between -4 and -1.5 (apart from R22 where log() reaches -0.5). For MP23, R22 and shock models, we could also test the effect of different densities (Ā Ā cm-3 in MP23, R22; Ā Ā cm-3 in shock models), finding small differences that only slightly impact the results shown in Sec.Ā 4, as we discuss in Sec.Ā 5. Here below we give a summary of the characteristics of all these models.
MP23 SSP models combine CLOUDY (Ferland etĀ al. 2017) with the revised version of the Bruzual & Charlot (2003) stellar population synthesis models introduced in Plat etĀ al. (2019; C&B models), spanning across several parameters (see their TableĀ 4), including a Kroupa (2001) and top-heavy (x030) initial mass functions (IMFs) with a mass range up to 100 and 300 Mā. C&B models accurately cover the evolution of O and B stars, including the Wolf-Rayet (WR) phase, whose contributions last a few Myr. C&B models also include hot post-Asymptotic Giant Branch (pAGB) stars, also known as hot low-mass evolved stars (HOLMES; StasiÅska etĀ al. 2008), which emit a strong UV continuum ionizing and exciting the surrounding medium at ages between 70 and 500 Myr (peak around 100 Myr).
@@
R22 made emission line predictions for Ā Mā IMBHs with CLOUDY, taking into account two extreme spectral energy distributions (SED) - a ādisk-plawā and a āqsosedā - for the incident radiation field (see R22 Sec.Ā 2.1) - and a starburst continuum including WR contribution, using the stellar population synthesis code BPASSĀ v2.0 (Stanway etĀ al. 2016) with a Kroupa IMF. To address different physical uncertainties, R22 models at different values of metallicities and ionization parameters include both coincident and non-coincident mixing of the IMBH and starlight excitation, and open and closed geometries. In particular, in this paper we show the 20Ā Myr instantaneous burst of star formation grids (in SSPs [OĀ III]Ā 5007/HĀ peaks at 20Ā Myr; Xiao etĀ al. 2018), with coincident mixing and closed geometry. The line ratios do not vary significantly in the non-coincident outputs, but can be slightly affected by the choice of geometry, as we comment in Sec.Ā 4.
AM19 shock models are made with Mappings-V (Sutherland & Dopita 2017), a plane-parallel geometry and the same prescriptions of the well-known Allen etĀ al. (2008) models, but spanning a broad range of gas-phase metallicities and densities. F24 models are made similarly with Mappings-V, but uniformly implemented empirical abundance patterns and dust depletion treatment as well as updated atomic data. Generally, two sets of models are taken into account: pure shocks and shocks+precursor. In this work, we consider only shocks+precursor models, since they also consider the gas entering the shock front. The grids we show span values of shock velocity (Ā km/s) and the magnetic field parameter (Ā Gauss), which are the main parameters regulating the shock ionization spectrum and the effective ionization parameter. We find that AM19 and F24 shock models behave similarly in the range of parameters considered, thus in Sec.Ā 4 we show only F24 grids.
@@ 4.1 Emission-line maps
Fig.Ā 5 shows the emission line maps of the extended MIR emission lines obtained after the point source subtraction: PfĀ Ā 7.49m, HuĀ Ā 12.37m, [NeĀ II]Ā Ā 12.81m, [SĀ III]Ā Ā 18.71m, [ArĀ III]Ā Ā 8.99m, [SĀ IV]Ā Ā 10.51m, [NeĀ III]Ā Ā 15.56m, [OĀ IV]Ā Ā 25.89m, and [NeĀ V]Ā Ā 14.32m, in order of ionization potential (see Tab.Ā 1). The maps of all the lines but [SĀ III]Ā and [OĀ IV]Ā (Channel 4; ā, Ā pc) are convolved with the spatial resolution at the [NeĀ III]Ā wavelength (ā, Ā pc) and rescaled to Channel 3 pixel scale (0.245ā). The FOV varies from 3.2ā3.7ā, to 4.0ā4.8ā, 5.2ā6.2ā and 6.6ā7.7ā for Channel 1 (Pf), 2 ([ArĀ III], [SĀ IV]), 3 (Hu, [NeĀ II], [NeĀ V], [NeĀ III]) and 4 ([SĀ III], [OĀ IV]), respectively. In the [SĀ III]Ā map panel we indicate the positions of SBSĀ 0335-052Ā E several star-forming regions as we did in Fig.Ā 1.
@@
We notice that the peak of [NeĀ III]Ā and [SĀ IV]Ā emission is close to the ULX position revealed by Prestwich etĀ al. (2013; see also ) with Chandra (magenta circle). However, there is no clear correspondence between the ULX X-ray emission and high ionization lines, since the [NeĀ V], [OĀ IV]Ā and UV emission peaks are shifted towards S7. We highlight that the ULX extracted spectrum is also characterized by [FeĀ II]Ā 5.34m emission, that is not shown in Fig.Ā 5 since it does not show clear extended emission (see Fig.Ā 12).
The region where we subtracted the point-source emission, located at the position of SSC1, shows either a reduced or tentative emission (grey regions with ) in the line maps. This can be clearly seen in the HuĀ and [NeĀ V]Ā maps. This could be due to a real lack of extended emission around SSC1 or to uncertainties in the PSF subtraction procedure. As shown in Fig.Ā 2 the point source SSC1 shows almost all the emission lines reported in Fig.Ā 5, apart from [NeĀ V]Ā and [ArĀ II]. The point source spectrum also shows clear [FeĀ II]Ā emission (see also Fig.Ā 12).
@@ 4.2 Emission-line ratio maps
Fig.Ā 6 shows [NeĀ III]/[NeĀ II], [NeĀ V]/[NeĀ II]Ā and [OĀ IV]/[NeĀ II]Ā line ratio extended emission maps with a color-coding chosen to enhance the peak regions to aid with the interpretation of the diagnostic diagrams. [NeĀ III]/[NeĀ II]Ā is known to be a good tracer of the ionization parameter log(), while [NeĀ V]/[NeĀ II]Ā and [OĀ IV]/[NeĀ III]Ā are found to be sensitive also to the hardness of the ionizing spectrum (Richardson etĀ al. 2022; Garofali etĀ al. 2024). Indeed, these line ratios do not show the same behavior. The peak of [NeĀ V]/[NeĀ II]Ā and [OĀ IV]/[NeĀ III]Ā maps are located around region S7, aligning with the peak in HĀ emission (see black contours). While [NeĀ III]/[NeĀ II]Ā is also enhanced in S7, it instead peaks at the position of S3-Pa, also bright in HĀ (see black contours). SSC3, SSC4, SSC5 and SSC6 are characterized by high ionization emission, but lower [NeĀ V]/[NeĀ II]Ā and [OĀ IV]/[NeĀ III]Ā line ratios than S7. As mentioned above, the extended region covering SSC1 (point source) and SSC2 does not have sufficient S/N in [NeĀ II], [NeĀ V]Ā and [OĀ IV]. Finally, S8 does not have either [NeĀ V]Ā or [OĀ IV]Ā emission but shows intermediate values of [NeĀ III]/[NeĀ II].
@@ 4.3 Ionized gas kinematics maps
Fig.Ā 7 shows the [SĀ IV], [NeĀ III]Ā and [NeĀ V]Ā velocity and intrinsic velocity dispersion maps. The velocity maps are centred on the systemic velocity of this galaxy according to previous optical studies (Ā km/s; Moiseev etĀ al. 2010)333 We centred the [NeĀ III]Ā velocity map -24Ā km/s below the systemic velocity Ā km/s found by Moiseev etĀ al. (2010). In particular, we find this specific systematic shift in HuĀ and [NeĀ II]Ā (Ch.Ā 3A), and [NeĀ III]Ā (Ch.Ā 3C) velocity maps also obtained for the Stage 3 products, so it does not depend on the point-source subtraction procedure explained in Sec.Ā A. We suspect that this shift is not connected to the physical conditions of the gas, as it corresponds to Ā half the wavelength step in ChannelĀ 3. Also, Pf, [ArĀ III]Ā and [OĀ IV], located in different channels and characterized by different I.P., show consistent velocity maps to [SĀ IV]Ā and [NeĀ V]., while the intrinsic velocity dispersion maps are obtained by subtracting in quadrature the instrumental broadening dependent on wavelength from the observed velocity dispersion (see Eq.Ā 1, Jones etĀ al. 2023).
The MIRI/MRS FOV is limited to the central Ā kpc2 of SBSĀ 0335-052Ā E and does not cover the entire ionized gas distribution traced in HĀ by previous studies (e.g., Moiseev etĀ al. 2010; Herenz etĀ al. 2023), with signs of outflowing gas extending up to Ā kpc (Herenz etĀ al. 2023; in the north-west direction, along the galaxy minor axis). Thus, despite Moiseev etĀ al. (2010) showing that on larger scales there is a disk-like north-west to south-east gradient, we only see the inner regions known to be characterized by perturbed kinematics (Izotov etĀ al. 2006; Moiseev etĀ al. 2010; Herenz etĀ al. 2017; 2023). In particular, [SĀ IV], [NeĀ III]Ā and [NeĀ V]Ā maps trace progressively higher ionization gas (Tab.Ā 1) and show a consistent velocity pattern, with a blue-shifted emission around the position of S3-Pa and in the north-west portion of the MIRI/MRS FOV (on the right of S7 and at the top of S8), and red-shifted emission from north to south and in the east region.
@@
The regions where [NeĀ III]Ā and [SĀ IV]Ā peaks probably have no [NeĀ V]Ā emission due to their lower ionization (they also partly show [NeĀ II]Ā emission; Fig.Ā 5) and are characterized by enhanced red-shifted velocities, possibly showing a flow of gas moving away from the [NeĀ V]Ā high ionization region. This picture is consistent with what found by Izotov etĀ al. (2006) with high-resolution () VLT/GIRAFFE ARGUS spectra, showing double-peaked HĀ profiles in the region north of SSCs 4-5 (see their Fig.Ā 7a and 8). In particular, they interpreted these features with the presence of an expanding shell of ionized gas at Ā km/s, while the HĀ broadening in SSCs 4-5 with higher dynamic activity than the surroundings.
@@ 4.4 Diagnostic diagrams
We highlight that [NeĀ V]/[NeĀ II]Ā versus [NeĀ III]/[NeĀ II]Ā represents the most reliable diagnostic diagram explored here, since these lines lie in Channel 3, where the attenuation law is flat, thus they are minimally impacted by PSF variations and dust correction. [SĀ IV]/[NeĀ II]Ā instead can be affected by dust reddening, while [OĀ IV]/[NeĀ III]Ā is the ratio most affected by the wavelength-dependent MIRI/MRS PSF variation correction, as well as residual fringing, given that [NeĀ III]Ā is located in ChannelĀ 3c and [OĀ IV]Ā in ChannelĀ 4c. For these reasons, we mainly focus this section on [NeĀ V]/[NeĀ II]Ā versus [NeĀ III]/[NeĀ II], moving the other two diagnostic diagrams in App.Ā B.
@@
It should be noted that for the MP23 SSP grids, the low metallicity 2.5 and 100Ā Myr SSPs show the highest line ratios for young and old stars, respectively. These two grids are the only ones (partially) able to cover the x- and y- ranges shown. However, we discarded the 100Ā Myr grid because old stars struggle to explain the observed number of ionizing photons observed in this system as we discuss in Sec.Ā 5.2.1. Also, we notice that the choice of the IMF does not affect the line ratios, but the choice of Ā Mā strongly enhances [NeĀ V]/[NeĀ II]. All the other MP23 grids and G24 SSPs and SXPs grids below 5Ā Myr fall outside the shown x- and y-ranges. R22 IMBHs grids with Ā Mā are also not shown because they cannot reproduce log([NeĀ III]/[NeĀ II]) higher than . The grids with Ā Mā instead can overlap to the observed line ratios at 8% AGN fraction, but at lower/higher fractions, they predict larger/lower [NeĀ III]/[NeĀ II].
@@
Another important takeaway from the top two panels of Fig.Ā 8 (same for Fig.Ā 13 and 14, App.Ā B) is that all the low-metallicity young SSPs grids struggle to reproduce the observed MIR emission line ratios of SBSĀ 0335-052Ā E. This suggests the need for a further ionization source. However, SXP models, including the contribution of X-ray binaries, show similar limitations. The only models that can reproduce SBSĀ 0335-052Ā E MIR [NeĀ V]/[NeĀ II], [NeĀ III]/[NeĀ II]Ā and [SĀ IV]/[NeĀ II]Ā line ratios are models including the contribution of an IMBH (%), spanning a range of log(). However, we highlight that these IMBH models still struggle to reproduce the highest [OĀ IV]/[NeĀ III]Ā line ratios (Fig.Ā 14). Literature BCDs [SĀ IV]/[NeĀ II]Ā vs [NeĀ III]/[NeĀ II]Ā observed with Spitzer follow the Groves etĀ al. 2008 empirical relation shown in Fig.Ā 13 (magenta line). This indicates that low-metallicity SSP, SXP and shock grids generally struggle to reproduce BCDsā lower ionization MIR lines than [NeĀ V].

### v2

@@ 3.4 Photoionization and shock models
SSP models with and without the ionizing output from X-ray binaries (XRBs) from Garofali etĀ al. (2024; G24 hereafter)222https://github.com/kgarofali/sxp-cloudy;
@@
Intermediate-massive black hole (IMBH) models from Richardson etĀ al. (2022; R22 hereafter)333https://github.com/crichardson17/richardson_2022;
@@
Shock models from Flury etĀ al. (2024; F24 hereafter)444https://zenodo.org/records/14140949 and the Mexican Million Models Shock Database555http://3mdb.astro.unam.mx:3686/, described in Alarie & Morisset (2019; AM19 hereafter).
For all these models, we considered only the two lowest metallicity grids available (i.e., 5% and 10% Zā models for G24, R22, F24, and 2% and 10% Zā for MP23), in accordance with SBSā0335-052Ā Eās metallicity range (Papaderos etĀ al. 2006; Nakajima etĀ al. 2024). Also, we took into account only bursts of star-formation with ages below 100Ā Myr (upper limit on age of SBS 0335-052Ā E) to match SBS 0335-052Ā E properties (Izotov etĀ al. 1997; Wofford etĀ al. 2021; Berg etĀ al. 2022; Mingozzi etĀ al. 2024, Martinez et al. in prep.). For all the grids, we consider the ionization parameter log() between -4 and -1.5 (apart from R22 where log() reaches -0.5). For MP23, R22 and shock models, we could also test the effect of different densities (Ā Ā cm-3 in MP23, R22; Ā Ā cm-3 in shock models), finding small differences that only slightly impact the results shown in Sec.Ā 4, as we discuss in Sec.Ā 5. In the following, we give a summary of the characteristics of all these models.
MP23 SSP models combine CLOUDY (Ferland etĀ al. 2017) with the revised version of the Bruzual & Charlot (2003) stellar population synthesis models introduced in Plat etĀ al. (2019; C&B models), spanning across several parameters (see their Tab.Ā 4), including a Kroupa (2001) and top-heavy (x030) initial mass functions (IMFs) with a mass range up to 100 and 300 Mā. C&B models accurately cover the evolution of O and B stars, including the Wolf-Rayet (WR) phase, whose contributions last a few Myr. C&B models also include hot post-Asymptotic Giant Branch (pAGB) stars, also known as hot low-mass evolved stars (HOLMES; StasiÅska etĀ al. 2008), which emit a strong UV continuum ionizing and exciting the surrounding medium at ages between 70 and 500 Myr (peak around 100 Myr).
@@
R22 made emission line predictions for Ā Mā IMBHs with CLOUDY, taking into account two extreme spectral energy distributions (SED) - a ādisk-plawā and a āqsosedā - for the incident radiation field (see R22 Sec.Ā 2.1) - and a starburst continuum including WR contribution, using the stellar population synthesis code BPASSĀ v2.0 (Stanway etĀ al. 2016) with a Kroupa IMF. To address different physical uncertainties, R22 models at different values of metallicities and ionization parameters include both coincident and non-coincident mixing of the IMBH and starlight excitation, and open and closed geometries. In particular, in this paper we show the 20Ā Myr instantaneous burst of star formation grids (in SSPs [OĀ III]Ā 5007/HĀ peaks at 20Ā Myr; Xiao etĀ al. 2018), with coincident mixing and different geometry. The line ratios do not vary significantly in the non-coincident outputs, but can be slightly affected by the choice of geometry, as we comment in Sec.Ā 4.
AM19 shock models are made with MAPPINGS-V (Sutherland & Dopita 2017), a plane-parallel geometry and the same prescriptions of the well-known Allen etĀ al. (2008) models, but spanning a broad range of gas-phase metallicities and densities. F24 models are made similarly with MAPPINGS-V, but uniformly implemented empirical abundance patterns and dust depletion treatment as well as updated atomic data. Generally, two sets of models are taken into account: pure shocks and shocks+precursor. In this work, we consider only shocks+precursor models, since they also consider the gas entering the shock front. The shown grids span values of shock velocity (Ā km/s) and the magnetic field parameter (Ā Gauss), which are the main parameters regulating the shock ionization spectrum and the effective ionization parameter. We find that AM19 and F24 shock models behave similarly in the range of parameters considered, thus in Sec.Ā 4 we show only F24 grids.
@@ 4.1 Emission-line maps
Fig.Ā 5 shows the emission line maps of the extended MIR emission lines obtained after the point source subtraction: PfĀ Ā 7.49m, HuĀ Ā 12.37m, [NeĀ II]Ā Ā 12.81m, [SĀ III]Ā Ā 18.71m, [ArĀ III]Ā Ā 8.99m, [SĀ IV]Ā Ā 10.51m, [NeĀ III]Ā Ā 15.56m, [OĀ IV]Ā Ā 25.89m, and [NeĀ V]Ā Ā 14.32m, in order of ionization potential (see Tab.Ā 1). The maps of all the lines but [SĀ III]Ā and [OĀ IV]Ā (Channel 4; ā, Ā pc) are convolved with the spatial resolution at the [NeĀ III]Ā wavelength (ā, Ā pc) and rescaled to Channel 3 pixel scale (0.245ā). The FOV varies from 3.2ā3.7ā, to 4.0ā4.8ā, 5.2ā6.2ā and 6.6ā7.7ā for Channel 1 (Pf), 2 ([ArĀ III], [SĀ IV]), 3 (Hu, [NeĀ II], [NeĀ V], [NeĀ III]) and 4 ([SĀ III], [OĀ IV]), respectively. In the [SĀ III]Ā map panel, we indicate the positions of several star-forming regions in SBSĀ 0335-052Ā E as we did in Fig.Ā 1.
@@
We notice that the peak of [NeĀ III]Ā and [SĀ IV]Ā emission is close to the ULX position revealed by Prestwich etĀ al. (2013; their TableĀ 3; see also ) with Chandra (magenta circle). However, there is no clear correspondence between the ULX X-ray emission and high ionization lines, since the [NeĀ V], [OĀ IV]Ā and UV emission peaks are shifted towards S7. We highlight that the ULX extracted spectrum is also characterized by [FeĀ II]Ā 5.34m emission, that is not shown in Fig.Ā 5 since it does not show clear extended emission (see Fig.Ā 12).
The region where we subtracted the point-source emission, located at the position of SSC1, shows either a reduced or tentative emission (gray regions with ) in the line maps. This can be clearly seen in the HuĀ and [NeĀ V]Ā maps. This could be due to a real lack of extended emission around SSC1 or to uncertainties in the PSF subtraction procedure. As shown in Fig.Ā 2 the point source SSC1 shows almost all the emission lines reported in Fig.Ā 5, apart from [NeĀ V]Ā and [ArĀ II]. Finally, we extracted a spectrum centered at the point source position in the residual cube (i.e., after point-source subtraction), summing all the spaxels within one PSF FWHM, confirming that the absence of [NeĀ V]Ā is not due to residual fringing in ChannelĀ 3 (see SecĀ 2). The point source spectrum also shows clear [FeĀ II]Ā emission (see also Fig.Ā 12).
@@ 4.2 Emission-line ratio maps
Fig.Ā 6 shows [NeĀ III]/[NeĀ II], [NeĀ V]/[NeĀ II]Ā and [OĀ IV]/[NeĀ III]Ā line ratio extended emission maps with a color-coding chosen to enhance the peak regions to aid with the interpretation of the diagnostic diagrams. [NeĀ III]/[NeĀ II]Ā is known to be a good tracer of the ionization parameter log(), while [NeĀ V]/[NeĀ II]Ā and [OĀ IV]/[NeĀ III]Ā are found to be sensitive also to the hardness of the ionizing spectrum (Richardson etĀ al. 2022; Garofali etĀ al. 2024). Indeed, these line ratios do not show the same behavior. The peak of [NeĀ V]/[NeĀ II]Ā and [OĀ IV]/[NeĀ III]Ā maps are located around region S7, aligning with the peak in HĀ emission (see black contours). While [NeĀ III]/[NeĀ II]Ā is also enhanced in S7, it instead peaks at the position of S3-Pa, also bright in HĀ (see black contours). SSC3, SSC4, SSC5 and SSC6 are characterized by high ionization emission, but lower [NeĀ V]/[NeĀ II]Ā and [OĀ IV]/[NeĀ III]Ā line ratios than S7. As mentioned above, the extended region covering SSC1 (point source) and SSC2 does not have sufficient S/N in [NeĀ II], [NeĀ V]Ā and [OĀ IV]. Finally, S8 does not have either [NeĀ V]Ā or [OĀ IV]Ā emission but shows intermediate values of [NeĀ III]/[NeĀ II]. Tab.Ā 2 reports the line ratios for the different regions discussed above.
@@ 4.3 Ionized gas kinematics maps
Fig.Ā 7 shows the [SĀ IV], [NeĀ III]Ā and [NeĀ V]Ā velocity and intrinsic velocity dispersion maps. The velocity maps are centred on the systemic velocity of this galaxy according to previous optical studies (Ā km/s; Moiseev etĀ al. 2010)666 We centred the [NeĀ III]Ā velocity map -24Ā km/s below the systemic velocity Ā km/s found by Moiseev etĀ al. (2010). In particular, we find this specific systematic shift in HuĀ and [NeĀ II]Ā (Ch.Ā 3A), and [NeĀ III]Ā (Ch.Ā 3C) velocity maps also obtained for the Stage 3 products, so it does not depend on the point-source subtraction procedure explained in Sec.Ā A. We suspect that this shift is not connected to the physical conditions of the gas, as it corresponds to Ā half the wavelength step in ChannelĀ 3. Also, Pf, [ArĀ III]Ā and [OĀ IV], located in different channels and characterized by different I.P., show consistent velocity maps to [SĀ IV]Ā and [NeĀ V]., while the intrinsic velocity dispersion maps are obtained by subtracting in quadrature the instrumental broadening dependent on wavelength from the observed velocity dispersion (see Eq.Ā 1, Jones etĀ al. 2023).
The MIRI/MRS FOV is limited to the central Ā kpc2 of SBSĀ 0335-052Ā E and does not cover the entire ionized gas distribution traced in HĀ by previous studies (e.g., Moiseev etĀ al. 2010; Herenz etĀ al. 2023), with signs of outflowing gas extending up to Ā kpc (Herenz etĀ al. 2023; in the north-west direction, along the galaxy minor axis). Thus, despite Moiseev etĀ al. (2010) showing that on larger scales there is a disk-like north-west to south-east gradient, we only see the inner regions known to be characterized by perturbed kinematics (Izotov etĀ al. 2006; Moiseev etĀ al. 2010; Herenz etĀ al. 2017; 2023). In particular, [SĀ IV], [NeĀ III]Ā and [NeĀ V]Ā maps trace progressively higher ionization gas (Tab.Ā 1) and show a consistent velocity pattern, with a blueshifted emission around the position of S3-Pa and in the north-west portion of the MIRI/MRS FOV (on the right of S7 and at the top of S8), and redshifted emission from north to south and in the east region.
@@
The regions where [NeĀ III]Ā and [SĀ IV]Ā peaks probably have no [NeĀ V]Ā emission due to their lower ionization (they also partly show [NeĀ II]Ā emission; Fig.Ā 5) and are characterized by enhanced redshifted velocities, possibly showing a flow of gas moving away from the [NeĀ V]Ā high ionization region. This picture is consistent with what found by Izotov etĀ al. (2006) with high-resolution () VLT/GIRAFFE ARGUS spectra, showing double-peaked HĀ profiles in the region north of SSCs 4-5 (see their Fig.Ā 7a and 8). In particular, they interpreted these features with the presence of an expanding shell of ionized gas at Ā km/s, while the HĀ broadening in SSCs 4-5 with higher dynamic activity than the surroundings.
@@ 4.4 Diagnostic diagrams
We highlight that [NeĀ V]/[NeĀ II]Ā versus [NeĀ III]/[NeĀ II]Ā represents the most reliable diagnostic diagram explored here, since these lines lie in Channel 3, where the attenuation law is flat, thus they are minimally impacted by PSF variations and dust correction. [SĀ IV]/[NeĀ II]Ā instead can be affected by dust reddening, while [OĀ IV]/[NeĀ III]Ā is the ratio most affected by the wavelength-dependent MIRI/MRS PSF variation correction, as well as residual fringing, given that [NeĀ III]Ā is located in ChannelĀ 3c and [OĀ IV]Ā in ChannelĀ 4c. This is also likely affecting the [OĀ IV]/[NeĀ III]Ā value estimated for the ULX extracted spectrum, that is Ā dex higher than the median value of the spatially resolved spaxels (see Fig.Ā 14, Tab.Ā 2). For these reasons, we mainly focus this section on [NeĀ V]/[NeĀ II]Ā versus [NeĀ III]/[NeĀ II], moving the other two diagnostic diagrams in App.Ā B.
@@
It should be noted that for the MP23 SSP grids, the low metallicity 2.5 and 100Ā Myr SSPs show the highest line ratios for young and old stars, respectively. These two grids are the only ones (partially) able to cover the x- and y- ranges shown. However, we discarded the 100Ā Myr grid because old stars struggle to explain the observed number of ionizing photons observed in this system, as we discuss in Sec.Ā 5.2.1. Also, we notice that the choice of the IMF does not affect the line ratios, but the choice of Ā Mā strongly enhances [NeĀ V]/[NeĀ II]. All the other MP23 grids and G24 SSPs and SXPs grids below 5Ā Myr fall outside the shown x- and y-ranges. R22 IMBHs grids with Ā Mā are also not shown because they cannot reproduce log([NeĀ III]/[NeĀ II]) higher than . The grids with Ā Mā instead can overlap to the observed line ratios at 8% AGN fraction, but at lower/higher fractions, they predict larger/lower [NeĀ III]/[NeĀ II].
@@
Another important takeaway from the top two panels of Fig.Ā 8 (same for Fig.Ā 13 and 14, App.Ā B) is that all the low-metallicity young SSPs grids struggle to reproduce the observed MIR emission line ratios of SBSĀ 0335-052Ā E. This suggests the need for a further ionization source. However, SXP models, including the contribution of X-ray binaries, show similar limitations. The only models that can reproduce SBSĀ 0335-052Ā E MIR [NeĀ V]/[NeĀ II], [NeĀ III]/[NeĀ II]Ā and [SĀ IV]/[NeĀ II]Ā line ratios are models including the contribution of an IMBH (%), spanning a range of log(). However, we highlight that these IMBH models still struggle to reproduce the highest [OĀ IV]/[NeĀ III]Ā line ratios (Fig.Ā 14). [SĀ IV]/[NeĀ II]Ā vs [NeĀ III]/[NeĀ II]Ā ratios observed with Spitzer in BCDs are consistent with the empirical relation from Groves etĀ al. 2008, shown as the magenta line in Fig.Ā 13. Since low-metallicity SSP, SXP, and shock models fail to reproduce this relation, they also struggle to match the lower-ionization MIR lines (below [NeĀ V]) typically observed in BCDs.